OpenStudy (jaylelile):

please, help me. 10. What are the domain, range, and asymptote of h(x) = 2x + 4? domain: {x | x > 0}; range: {y | y is a real number}; asymptote: y = 0 domain: {x | x > –4}; range: {y | y is a real number}; asymptote: y = –4 domain: {x | x is a real number}; range: {y | y > 0}; asymptote: y = 0 domain: {x | x is a real number}; range: {y | y > 0}; asymptote: y = –4

OpenStudy (jaylelile):

8. What is the domain of f(x) = 5x – 7? {x | x > –7} {x | x < –7} {x | x > 0} {x | x is a real number}

OpenStudy (jaylelile):

4. The graph of g(x) is the result of translating the graph of f(x) = 3x six units to the right. What is the equation of g(x)? g(x) = 3x – 6 g(x) = 3x + 6 g(x) = 3x – 6 g(x) = 3x + 6

OpenStudy (jaylelile):

2. What are the domain, range, and asymptote of h(x) = 6x – 4? domain: {x | x is a real number}; range: {y | y > 4}; asymptote: y = 4 domain: {x | x is a real number}; range: {y | y > –4}; asymptote: y = –4 domain: {x | x > –4}; range: {y | y is a real number}; asymptote: y = 4 domain: {x | x > 4}; range: {y | y is a real number}; asymptote: y = –4

OpenStudy (jaylelile):

@LynFran ??

OpenStudy (misssunshinexxoxo):

Could you kindly post 1 question at a time?

OpenStudy (lynfran):

for question8. x is a real #

OpenStudy (jaylelile):

is the +6 an exponant?

OpenStudy (lynfran):

no

OpenStudy (jaylelile):

ok

OpenStudy (anonymous):

A line that a curve approaches, as it heads towards infinity is an asymptote.... This h(x) = 2x + 4 is just a straight line/ linear equation. It has no asymptote

OpenStudy (lynfran):

im not sure about the asymptote part but..for question 10...x>-4 and y is a real # so im thinking it A or B

OpenStudy (lynfran):

i think its B though

OpenStudy (lynfran):

for question #2, y is a real number..i think this 1 is C